### v1

# Fast and Globally Consistent Normal Orientation based on the Winding Number Normal Consistency

Abstract
Estimating a consistently oriented normal vector field for an unoriented point cloud enables a number of important downstream applications in computer graphics. While normal estimation for a small patch of points can be done with simple techniques like principal component analysis (PCA), orienting these normals to be globally consistent has been a notoriously difficult problem. Some recent methods exploit various properties of the winding number formula to achieve global consistency with state-of-the-art performance. Despite their exciting progress, these algorithms either have high space/time complexity, or do not produce accurate and consistently oriented normals for imperfect data. In this paper, we derive a novel property from the winding number formula to tackle this problem: the normal consistency property of the winding number formula. We refer to this property as the winding number normal consistency (WNNC). The derived property is based on the simple observation that the normals (negative gradients) sampled from the winding number field should be codirectional to the normals used to compute the winding number field. We further propose to turn the WNNC property into a normal update formula, which leads to an embarrassingly simple yet effective iterative algorithm that allows fast and high-quality convergence to a globally consistent normal vector field. Furthermore, our proposed algorithm only involves repeatedly evaluating the winding number formula and its derivatives, which can be accelerated and parallelized using treecode-based approximation algorithms due to their special structures. Exploiting this fact, we implement a GPU-accelerated treecode-based solver. Our GPU (and even CPU) implementation can be significantly faster than the recent state-of-the-art methods for normal orientation from raw points.
[…]
Figure 1: Given a raw point cloud, our method iteratively updates its normals to be globally consistently oriented (points rendered as spheres colored by normals). Our iterative algorithm produces overall consistent normals even at the first iteration, and fully converges in around 40 iterations. With our treecode-accelerated implementation running on GPU (RTX 3090), the algorithm obtains high-fidelity normals in $\sim$1 second for 50,000 points, and $\sim$30 seconds for a point cloud with half a million points.
1 Introduction
Point clouds are one of most widely used 3D representations in computer graphics applications such as geometric modeling and rendering. In most cases, consistently oriented normal vectors associated to points are necessary, e.g., for surface reconstruction [24, 23, 25, 32]. While locally estimating normals for a small patch of points can be done with simple techniques such as principal component analysis (PCA), obtaining a globally consistent orientation remains a notoriously difficult problem. Older methods mostly take a propagation-based approach by first estimating normals locally and then propagating the orientation globally [18, 42, 27, 37, 34]. These methods are generally fast but less robust to noise, thin structures and sharp edges where propagation easily fails. Recent developments focus more on the global consistency by introducing new optimization formulations that involve the whole point cloud [19, 31, 44]. Despite achieving the state-of-the-art performance, these global methods often suffer from high space/time complexity, preventing their practicability.
An exciting and inspiring tread in this field is the exploration of various properties of the winding number formula (see Theorem 6). From the second line in Eq. (5) of the winding number formula, Lin et al. [31] observed that the winding number field computed using correct normals should evaluate to $1/2$ at the input points. Lin et al. [31] further turned this observation into a practical linear system, from which normals can be solved using conjugate gradients [17]. Complementarily, Xu et al. [44] use the first and the third line in Eq. (5) to define an regularization term such that the winding number evaluates to either 00 or $1$ almost everywhere. By turning the winding number formula to optimization objectives, these two methods above successfully achieve state-of-the-art performance in estimating consistently oriented normals. However, both methods suffer from high computational complexity. The linear system in PGR [31] has size $N\times N$ and is dense. Running PGR [31] with 40,000 points consumes about 8GB GPU memory. The high memory usage prohibits it from scaling to large point clouds. On the other hand, GCNO [44] has high time complexity due to its objective function being non-linear. Solving a 10,000-point model using GCNO takes about one hour.
While it seems that PGR [31] and GCNO [44] have exhausted all possibilities with the winding number formula by using all three lines in Eq. (5) as constraints, we further exploit the winding number formula by exploring its higher-order properties. To be more specific, we propose to utilize the fact that, when given correctly oriented normals to evaluate the winding number formula, the normals (negative gradients) of the formula should be codirectional to the given normals. We refer to this property as the winding number normal consistency (WNNC). An illustration of WNNC is given in Fig. 2. Based on the WNNC, we design an iterative algorithm for updating normals that allows fast and high-quality convergence to a globally consistent normal vector field. Furthermore, the proposed algorithm only involves repeatedly evaluating the winding number formula and its derivatives. Due to their special structures, these operations can be accelerated using a treecode-based algorithm [4] and parallelized on GPU. Lacking an open-source implementation of the treecode algorithm that fully supports our needs (GPU-based and support for the winding number kernel derivative functions), we implement our custom treecode-based solver in CUDA. As shown in Table 1, our implementation significantly outperforms recent state-of-the-art methods in terms of space/time complexity.
We summarize our contributions as follows:
  • We derive the winding number normal consistency (WNNC) property from the winding number formula to provide a new mathematical formulation for estimating consistently oriented normals from raw points.
  • We design an iterative algorithm based on the WNNC property, which achieves fast and high-quality convergence to a globally consistent normal vector field.
  • We implement a treecode-based acceleration algorithm in CUDA, achieving substantial improvements over prior methods in terms of computational costs.
NormalsWinding number fieldNegativegradientsAligned ✓Not aligned ✗
Figure 2: Illustration of the winding number normal consistency (WNNC) property. From left to right, we start with an input point cloud with normals, evaluate the winding number formula, and finally compute the negative field gradients at the input point positions. If the normals at the start are correctly oriented as shown in the first row, then the intermediate field is exactly the winding number field of the underlying shape. Consequently, the negative gradients at these point positions would be aligned with the input normals. We refer to this property as the winding number normal consistency (WNNC) property. Note that random normals do not satisfy the WNNC property (second row).
Table 1: A brief comparison of time and memory consumption for the state-of-the-art methods GCNO [44], PGR [31], Dipole [34] and iPSR [19]. The results are obtained from running an Armadillo model [28] with uniform sampling on a desktop with an Intel i9-10900X CPU (3.70GHz, 10 cores) and an NVIDIA RTX 3090 GPU. Note that the cells colored in gray are estimates based on complexity, because GCNO and PGR are too costly to run on large-scale point clouds.
Method $10^{4}$ points $10^{5}$ points $10^{6}$ points
Time Mem. Time Mem. Time Mem.
GCNO (CPU) 1.1h $<$1G $>$1day 1.7G weeks $>$10G
PGR (GPU) 8s 1G $>$10min $>$30G $>$1day $>$3000G
Dipole (GPU) 17s 1G 90s 2.3G 230s 7.5G
iPSR (CPU) 18s $<$1G 50s $<$1G 380s 2.6G
Ours (CPU) 3s $<$1G 15s $<$1G 294s 1G
Ours (GPU) 2s $<$1G 6s $<$1G 86s 1.5G
2 Related Work
The research into estimating consistent normal orientation has a long history, dating back to decades ago. In this review, we mainly introduce methods that focus on the global consistency of normal orientation. Following prior work, we categorize existing methods into three types: propagation-based, volumetric and deep learning-based. We remark that these types are not completely mutually exclusive, and we introduce existing work based on their most prominent contribution.
2.1 Propagation-based Methods
Propagation-based methods mostly follow a two-step approach: They first estimate local normals with undetermined orientations, and then propagate the orientation from a seed point or local patch through its neighbors to the whole point cloud.
Local normal estimation is usually carried out by fitting a primitive geometry to local samples. Popular choices for this step include local tangent planes obtained from Principal Component Analysis (PCA) [18], implicit quadric surfaces [42, 43], polynomials obtained by moving least squares [29], osculating jets [6, 34]. These techniques are often efficient and well-established.
Once unoriented normals are available, it remains to determine a sign for each of them. The general practice is to propagate local orientations through neighbors to achieve global consistency. Hoppe et al. [18] build a graph for input points whose edge costs are given by a measure of orientation consistency. Seeking a globally consistent orientation is then formulated as a graph optimization problem. Hoppe et al. [18] use the minimum spanning tree (MST) as a greedy solution, which boils down to propagating the orientation along the MST. This simple greedy approach is efficient but lacks robustness. Many follow-up methods improve one or more aspects of Hoppe et al. [18]. To be robust to sharp features, Xie et al. [42] start from multiple seed points with an improved flipping criterion to orient the whole point cloud. König and Gumhold [27] further introduced Hermite curves into the flipping criterion to reduce the possibility of failures. Seversky et al. [37] use harmonic functions on the point cloud to improve the constructed MST. Going beyond MST-based approaches, other propagation schemes have been proposed as well. Jakob et al. [22] build a graph of points and equate a graph-edge collapse propagation with as an orientation step, and implement a fast parallel algorithm that collapse edges in a greedy manner. Instead of propagating orientations on the surface, Xie et al. [43] first use locally oriented surfaces to determine locally oriented regions, and then propagate region orientations through a voting scheme. Recently, Metzer et al. [34] propose to use neural networks to estimate consistently orientated normals for local patches, and then use dipole-based propagation to achieve global consistency.
These propagation-based methods are generally easily formulated and efficient. However, they often fail to handle difficulties such as noise, sharp edges or holes. These methods are also very sensitive to a number of factors, e.g., neighborhood size for propagation, flipping criterion, graph weight selection, etc. Thus, the applicability of these methods are limited, especially when dealing with complex geometries or imperfections.
2.2 Volumetric Methods
Instead of directly estimating normals for point clouds, volumetric methods adopt an indirect approach where the whole volume is partitioned into inside/outside regions. Normal orientations can be then be determined as the direction pointing from inside to outside.
Early works make use of space partitioning structures to identify inside/outside regions. Dey and Goswami [10] prove that, under certain assumptions, Delaunay balls of the Voronoi diagram can well separated into outer balls and inner balls to give a desired partitioning. Chen et al. [7] uses an octree, called binary orientation tree, that gradually tags its node corners and inside/outside as it grows.
Later mainstream volumetric methods mostly seek an implicit field $f(x)$, and define the inside region as $\{x:f(x)<c\}$ or $\{x:f(x)>c\}$. These methods usually regularize the desired implicit field with extra constraints to ensure global consistency, or directly solve for commonly used implicit fields (e.g., the signed distance field and the winding number field).
For applying additional constraints, Alliez et al. [1] make use of unoriented normals estimated from Voronoi cells, and solve for an implicit field that minimizes an anisotropic Dirichlet energy measuring the unoriented alignment between Voronoi-based normals and the implicit field gradient. Walder et al. [40] apply value and gradient maximization together with smoothness regularization to an implicit function represented by radial basis functions (RBFs). Also based on RBFs, Huang et al. [20] adopt triharmonic kernels and Duchon’s energy [11] and reported high robustness to point sparsity and non-uniform sampling.
Signed distance fields (SDFs) are a popular choice of implicit fields. Mello et al. [33] compute an approximate SDF on a tetrahedral grid. Zhao et al. [46] solves the Eikonal equation to obtain the SDF on a rectangular grid. Gropp et al. [15] turns the Eikonal equation as a loss function supervising a neural network representation of the SDF.
Most closely related to our work are methods based on the winding number field [3, 32]. Simply speaking, the winding number field of an object equals $1$ on the inside and 00 on the outside. Since computing the winding number field requires known normals, several methods derive properties of the winding number field and use them in turn as constraints to obtain normals. Observing the fact that the winding number field is piecewise constant (i.e., having zero Dirichlet energy), Takayama et al. [38] solve for an orientation that minimizes the Dirichlet energy of the field computed with the winding number formula. Recently, Lin et al. [31] derived a linear system formulation enforcing the field to equal $1/2$ on input samples, from which normals can be efficiently solved. Xu et al. [44] designed a double-well loss function encouraging the field to evaluate to either 00 or $1$ almost everywhere.
Compared with propagation-based methods, most volumetric methods take a more “global” view, in the sense that their optimization objectives considers all input points throughout¸ the whole algorithm execution process. As a result, these methods generally achieve much better global consistency but at a notably higher computational cost.
2.3 Deep Learning Methods
In the past decade, deep learning has stormed over nearly all scientific fields, with normal estimation being no exception. The powerful fitting ability of neural networks allows direct inference of normals given an input point cloud. PCPNet [16] and Nesti-Net [5] use neural networks to process multi-scale point features and directly infer point normals. There are also volumetric learning-based methods that infer the implicit rather than the normals [12, 41]. Different from these straight-forward regression approaches, AdaFit [48] predicts point offsets prior to normal estimation to reduce the number of outliers. NeAF [30] further extends this idea by learning a neural angle field that offsets any given normal to the ground-truth one.
Benefiting from their ability to learn from data, deep learning methods generally exhibit better robustness to data imperfections such as noise or missing data. However, their performance strongly depends on training data and model capacity. Since current datasets are limited in quantity, complexity and variety compared to real-life objects, deep learning methods still suffer from generalizability issues.
2.4 Other Consistent Normal Estimation Methods
A few methods do not belong to the aforementioned categories. Shape as Points (SAP) [36] derives a differentiable version of Poisson Surface Reconstruction (PSR) [24], and uses it to update normals such that the reconstructed mesh has a low error with the input points. Since the differentiable PSR [36] is performed in the frequency domain with fast Fourier transform, its reconstruction accuracy is limited by grid resolution. Iterative Poisson Surface Reconstruction (iPSR) [19] starts with applying Screened Poisson Surface Reconstruction (SPSR) [23] to points with arbitrary normals, and iteratively resample normals from the mesh reconstructed in the previous iteration. Hou et al. [19] show that this simple iterative procedure can update normals into a consistent orientation in only 5$\sim$30 of steps.
3 Preliminaries
Our method builds on the winding number formula [3] (also referred to as the Gauss formula in potential theory [32]), and is closely related to PGR [31]. We begin with a brief review of them.
Theorem 1 (Winding Number).
Let $\Omega\subset\mathbb{R}^{3}$ be an open and bounded region. Suppose its boundary $\partial\Omega$ is smooth and $n(y)$ is the outward unit normal at $y\in\partial\Omega$. Then for any $y=(y_{1},y_{2},y_{3})\in\mathbb{R}^{3}$,
$F(y)$ $=$ $\int_{\partial\Omega}((\nabla\Phi)(y-x))\cdot n(x)\,\mathrm{d}S(x)$ (1)
$=$ $\left\{\begin{array}[]{lll}0&,&y\in\bar{\Omega}^{c}\\ 1/2&,&y\in\partial\Omega\\ 1&,&y\in\Omega\end{array}\right.$ (5)
is the indicator function of $\Omega$. Here, $\Phi(y)$ is the fundamental solution to the $3$-dimensional Laplace equation:
$\Phi(y)=\frac{1}{4\pi|y|},\quad\nabla\Phi(y)=-\frac{y}{4\pi|y|^{3}}.$ (6)
With a known normal vector field $n(x)$, the indicator function can be directly evaluated by Theorem 6. To numerically evaluate the winding number given a point set $\mathcal{P}=\{x_{j}\}_{j=1}^{N_{\mathcal{P}}}$ that samples the surface $\partial\Omega$, Lin et al. [31] adopt the following simple discretization:
$F(y)$ $=$ $\int_{\partial\Omega}((\nabla\Phi)(y-x))\cdot n(x)\,\mathrm{d}S(x)$ (7)
$\approx$ $\sum_{j=1}^{N_{\mathcal{P}}}\underbrace{\frac{-(y-x_{j})}{4\pi|y-x_{j}|^{3}}}_{=((\nabla\Phi)(y-x_{j}))\in\mathbb{R}^{3}}\cdot\underbrace{n(x_{j})\,\sigma(x_{j})}_{=\mu_{j}\in\mathbb{R}^{3}}$
$=$ $\sum_{j=1}^{N_{\mathcal{P}}}((\nabla\Phi)(y-x_{j}))\cdot\mu_{j}.$ (8)
Here, $\sigma(x_{j})$ denotes the local surface element area, $n(x_{j})$ denotes the unit vector and $\mu_{j}$ denotes their product, i.e., the oriented surface element at $x_{j}$. Lin et al. [31] further noticed that (5) in fact defines a linear equation in $\mu_{j}$:
$\sum_{j=1}^{N_{\mathcal{P}}}(\nabla\Phi)(x_{i}-x_{j})\cdot\mu_{j}=1/2,\ \text{for all }x_{i}\in\mathcal{P}.$ (9)
Lin et al. [31] then rewrite Eq. (9) into a regularized symmetric system, which can then be efficiently solved with conjugate gradients [17]. Finally, $\mu_{j}$ are inserted back into Eq. (8) to compute the winding number field.
Our method also applies the discretization in Eq. (8) as the representation of an indicator function, and we also utilize the constraints Eq. (9) for obtaining the $\mu_{j}$’s. In addition, we propose an iterative scheme which incorporates the WNNC that is absent from previous work, together with a GPU-based treecode-accelerated implementation to improve the efficiency of our algorithm, detailed in Section 4.
Gradient stepWNNC updateRescaleNormalizeRaw pointsOne iterationFinal output
Figure 3: The overall pipeline of our iterative algorithm. Each iteration consists of three steps: a gradient step w.r.t. a quadratic energy function, a normal update step based on the WNNC property and a rescaling step to ensure numerical stability. Our algorithm can generally produce an overall correct orientation at the first iteration, and gradually converges to details in a few dozens of iterations.
4 Method
4.1 Notations
Throughout the paper, we use $\mathcal{P}=\{x_{j}\}_{j=1}^{N_{\mathcal{P}}}$ to denote the unoriented input points, which are assumed to distribute on or near the surface of a 3D shape. Our goal is to estimate normals [^1] Strictly speaking, $\mu_{j}$ are oriented surface elements with their lengths being surface element areas. However, in most cases it is only the orientation that matters. We thus use the word normals to refer to $\mu_{j}$ for simplicity. $\mu_{j}$ that are consistently oriented. For the remaining part of this paper, we only use the discretized winding number formula Eq. (9). We also rewrite it as $F(y;\mu)$ to highlight the dependency of $F$ on $\mu$. Furthermore, we consider $\mu$ as unknowns. Note that $F(y;\mu)$ does not represent the actual winding number unless $\mu$ represents the actual normal vector field, but we call $F(y;\mu)$ the winding number formula nonetheless for convenience. The rest of this paper is devoted to finding a solution $\mu$ that represents consistently oriented normals.
4.2 Winding Number Normal Consistency
In this section, we introduce our core contribution, the winding number normal consistency (WNNC). The WNNC property is based on a simple observation: If the given normals $\mu_{j}$ are consistently outward pointing, i.e., satisfying the assumptions in Theorem 6, then the winding number field $F(y;\mu)$ computed using these normals should be the indicator field of the corresponding shape. Consequently, the negative gradient of $F(y;\mu)$ at $x_{i}$ should point to the same direction as $\mu_{i}$. Fig. 2 illustrates two cases, where one starts with (possibly incorrectly oriented) normals $\mu$, evaluates $F(y;\mu)$ and then resamples its negative gradient at each $x_{i}$. Note that if the original normals $\mu$ are correctly oriented, the resampled ones would still be aligned with them. The WNNC property can be expressed mathematically as follows: If $\mu_{j}$ are consistently outward pointing, then
$\mu_{i}=\lambda_{i}(-\nabla F(x_{i};\mu))\quad\textrm{for some}\quad\lambda_{i}>0.$ (10)
Fortunately, it is not necessary to compute $F(y;\mu)$ first and then its gradient. The discretized expression of $\nabla F(y;\mu)$ can be directly derived from Eq. (8):
$\nabla F(y;\mu)$ $=$ $\nabla\Big{(}\sum_{j}((\nabla\Phi)(y-x_{j}))\cdot\mu_{j}\Big{)}$ (11)
$=$ $\sum_{j}((H\Phi)(y-x_{j}))\cdot\mu_{j},$
where $H\Phi$ is the Hessian of $\Phi$:
$H\Phi(y)=-\frac{I_{3}}{4\pi|y|^{3}}+\frac{3}{4\pi|y|^{5}}\left[\begin{matrix}y_{1}\\ y_{2}\\ y_{3}\end{matrix}\right][y_{1}\ y_{2}\ y_{3}].$ (12)
To simplify notations, we define the mapping $\mu\mapsto\hat{\mu}$, where $\hat{\mu}_{i}:=-\nabla F(x_{i};\mu)$, as $\hat{\mu}=G(\mu)$. In other words, $G$ takes normals $\mu$ as input, and evaluates the negative gradients of the winding number field at the same point locations. Note that the expression of $G$ depends only on the point locations $x_{j}\in\mathcal{P}$.
Finally we can conclude the WNNC property in a simple mathematical form: If $\mu_{j}$ are consistently outward pointing, then
$\mu_{i}=\lambda_{i}(-\nabla F(y_{i};\mu))=\lambda_{i}G(\mu)_{i}\quad\textrm{for some}\quad\lambda_{i}>0.$ (13)
4.3 WNNC-based Iterative Algorithm
In this section, we describe our algorithm that utilizes the WNNC property Eq. (13) to achieve fast and high-quality convergence. At first sight, one may think it is natural to define an objective function to minimize the discrepancy between the directions of $\mu$ and $G(\mu)$, e.g., using the cosine similarity between them. However, this would lead to a non-linear problem that may be complicated to solve. We observe that the WNNC condition Eq. (13) seems to hint $\mu$ is the fixed point of the mapping $\mu\mapsto G(\mu)$ (up to per-point scaling). Inspired by fixed-point iteration algorithm in classical numerical analysis, we attempt to adopt an iterative scheme for solving normals using the WNNC property, where we update $\mu$ using $\mu\mapsto G(\mu)$. We call $\mu\mapsto G(\mu)$ a WNNC update.
Despite the similarity between WNNC updates and fixed-point iterations, merely applying the fixed-point iteration algorithm is infeasible due to the following reasons. (a) A trivial solution $\mu=0$ is clearly a fixed point of $\mu\mapsto G(\mu)$. (b) Even if we avoid $\mu=0$, there is no guarantee of convergence when using WNNC updates. In practice we also find the WNNC update alone is insufficient for convergence. (c) The WNNC property involves only the normal directions, but says nothing about the scale changes $\lambda_{i}$. Thus, $\mu$ is not a fixed point of $\mu\mapsto G(\mu)$ in a traditional sense. Furthermore, we empirically found $\lambda_{i}$ to range from $10^{4}$ to $10^{10}$ for typical datasets, which would blow up numerically in only a few iterations.
Since WNNC updates alone are insufficient, we propose to incorporate the constraints Eq. (9) from PGR [31] as an additional objective function:
$E=\sum_{i}\Big{|}\sum_{j=1}^{N_{\mathcal{P}}}(\nabla\Phi)(x_{i}-x_{j})\cdot\mu_{j}-1/2\Big{|}^{2},$ (14)
Again, to simplify notations, we define the mapping $\mu\mapsto s$, where $s_{i}:=F(x_{i};\mu)$, as $s=A(\mu)$. In other words, $\mu\mapsto s=A(\mu)$ evaluates the winding number field at each $x_{i}$ with given normals $\mu$. In this way, we can rewrite $E$ as
$E(\mu;A,b)=\|A(\mu)-b\|^{2},\ \textrm{with}\ b=(1/2,\cdots,1/2)^{T}.$ (15)
Our iterative algorithm aims to update normals using both WNNC updates $\mu\mapsto G(\mu)$ and $E(\mu;A,b)$. More specifically, our algorithm starts with an all-zero initial vector field $\mu=0$. During each iteration, we first take a gradient step w.r.t. $E(\mu;A,b)$, and then update the normals using an WNNC update. To solve the scale change problem during WNNC updates mentioned above, we rescale each $\mu_{i}$ to have the same length after each WNNC update. The algorithm is summarized as Algorithm 1, and Fig. 3 presents a graphical overview of our iterative pipeline.
Data: Unoriented point set $\mathcal{P}=\{x_{j}\}$ with $N_{\mathcal{P}}$ points.
Result: Normals $\mu$.
$\mu\leftarrow 0\in\mathbb{R}^{N\times 3}$;
for $i\leftarrow 1$ to $n$ do
       $\mu\leftarrow{\rm grad\_step}(\mu;A,b)$;
        // as in Alg. 2
       $\hat{\mu}\leftarrow{G(\mu)}$;
        // WNNC update
       $\mu\leftarrow{\rm rescale}(\mu,\hat{\mu})$;
        // $\mu_{i}\leftarrow\hat{\mu}_{i}|\mu_{i}|/|\hat{\mu}_{i}|$
end for
Algorithm 1 WNNC-based iterative algorithm (no smoothing)
Note that $\mu\mapsto A(\mu)$ is a linear operator, and thus $E(\mu;A,b)$ is a quadratic form. A single gradient step of $\mu$ w.r.t. $E(\mu;A,b)$ is presented in Algorithm 2. This can be derived with basic linear algebra.
Data: Matrix $A$ and vectors $\mu,b$.
Result: Updated $\mu$ after a single gradient step w.r.t. $\|A\mu-b\|^{2}$.
$r\leftarrow A^{T}b-A^{T}A\mu$;
$\alpha\leftarrow r^{T}r/r^{T}A^{T}Ar^{k}$;
$\mu\leftarrow\mu+\alpha r$;
Algorithm 2 Single-step gradient update grad_step($\mu;A,b$).
4.4 Resolving Singularities
The steps in Algorithm 1 are none other than evaluating expressions of the form $(\nabla\Phi)(x_{i}-x_{j})$ and $(H\Phi)(x_{i}-x_{j})$. However, both expressions become singular when $x_{i}$ approaches $x_{j}$. Using their raw forms inevitably brings numerical instability. Following previous work [32], we set their values to 00 if $\|x_{i}-x_{j}\|<w$, where $w>0$ is a constant called the smoothing width. This change affects the operators $A$ and $G$. We denote their modified versions by $A_{w}$ and $G_{w}$, respectively. As reported in previous work [32, 31], a small $w$ retains more details but is less robust to noise and other imperfections, while a large $w$ leads to an overly smooth winding number field. In practice, we find that using a fixed width as in PGR [31] either leads to oversmoothing or cannot ensure global consistency. We thus propose a width scheduling scheme, where we start with a large $w=w_{2}$ and linearly decreases it to a small $w=w_{1}$ as the algorithms proceeds. In practice, This allows the iterative algorithm to remain stable during early iterations, and to enhance details in late iterations. The final algorithm is described in Algorithm 3.
Data: Unoriented point set $\mathcal{P}=\{x_{i}\}$ with $N_{\mathcal{P}}$ points, $w_{1}\leq w_{2}$.
Result: Normals $\mu$.
$\mu\leftarrow 0\in\mathbb{R}^{N\times 3}$;
for $i\leftarrow 1$ to $n$ do
       $w=w_{2}\frac{n-i}{n-1}+w_{1}\frac{i-1}{n-1}$;
       $\mu\leftarrow{\rm grad\_step}(\mu;A_{w},b)$;
        // as in Alg. 2
       $\hat{\mu}\leftarrow{G_{w}(\mu)}$;
        // WNNC update
       $\mu\leftarrow{\rm rescale}(\mu,\hat{\mu})$;
        // $\mu_{i}\leftarrow\hat{\mu}_{i}|\mu_{i}|/|\hat{\mu}_{i}|$
end for
Algorithm 3 WNNC-based iterative algorithm (with smoothing)
From now on we assume the smoothing width is always used, but for convenience we may not make it explicit in notations.
(a)(b)(c)(d)SourcesTargetRepresentatives
Figure 4: Illustration of the treecode-based approximation algorithm [4]. (a) We wish to sum up interactions between multiple source points and the target point. (b) The space is partitioned by an octree hosting the source points, and an representative point is computed in each node (not shown here). (c) Based on the proximity between the target point and the octree nodes, representative points are chosen. Attributes of the source points are aggregated to their corresponding representative points. (d) For far-field interactions, representative points are used instead of the original source points.
4.5 Treecode-based Acceleration on GPU
As described in Algorithm 2 and Algorithm 3, our iterative scheme mainly involves evaluating the three linear operators $\mu\mapsto A(\mu)$, $p\mapsto A^{T}(p)$ and $\mu\mapsto G(\mu)$. A naive method for evaluating each of these operator has time complexity $O(N_{\mathcal{P}}^{2})$. The computational time becomes prohibitively long as the number of input points grows.
The idea for accelerating these operators is based on the following observation: Each operator can be seen as summing up the interactions between $x_{i}$ (targets) and $x_{j}$ (sources) through kernel functions $\nabla\Phi(x_{i}-x_{j})$ or $\nabla H(x_{i}-x_{j})$, and since both kernels vanish sufficiently fast as $|x_{i}-x_{j}|$ increases, far-field interactions of a group of neighboring sources can be well approximated by a new representative source (see Fig. 4).
To build representative sources, we first build an octree $\mathcal{T}$ to spatially partition the input point set. The partitioning stops if the each leaf node contains only one point or if the user-specified maximum depth $D$ is reached. Then, at each iteration, each octree node accumulates point attributes (such as $\mu_{i}$) from the points it contains and select a representative point as a weighted average of all points in the node. The representative point of each node is determined as follows. Let $B$ be some node of the octree and let $\mathcal{P}_{B}$ be the subset of input points that lie in $B$. Let $\nu_{j}$ be an attribute vector of $x_{j}$ that we wish to propagate to $x_{i}$ through an interaction (e.g., $\nu_{j}=\mu_{j}$ when evaluating $\mu\mapsto A(\mu)$ or $\mu\mapsto G(\mu)$, and $\nu_{j}=s_{j}$ when evaluating $s\mapsto A^{T}s$). The $\nu$-representative of the node $B$ is located at
$x_{B,\nu}=\sum_{x_{i}\in\mathcal{P}_{B}}\frac{|\nu_{i}|}{\sum_{x_{j}\in\mathcal{P}_{B}}|\nu_{j}|}x_{i},$ (16)
associated with the representative attribute vector
$\nu_{B}=\sum_{x_{i}\in\mathcal{P}_{B}}\nu_{i}.$ (17)
With the $\nu$-representatives already obtained, the operators above can be evaluated by simple tree traversal. Algorithm 4 describes the process in detail using $\mu\mapsto s=A(\mu)$ as an example.
Data: Query point $x_{i}$; normals at the current iteration $\mu=(\mu_{i})\in\mathbb{R}^{N_{\mathcal{P}}\times 3}$; octree $\mathcal{T}$ with $\mu$-representatives already computed; the starting tree node $B$.
Result: An output scalar $s_{i}$ associated to $x_{i}\in\mathcal{P}$.
$s_{i}\leftarrow 0$;
if $|x_{i}-x_{B,\mu}|>c\cdot({\rm the\ width\ of\ }B)$ then
       $s_{i}\leftarrow s_{i}+(\nabla\Phi)(x_{i}-x_{B,\mu})\cdot\mu_{B}$;
        // modified by $w$
else if $B$ is not a leaf then
       for each child node $B_{\rm c}$ of $B$ do
             $s_{i}\leftarrow s_{i}+{\rm apply\_A}(x_{i},\mu,B_{\rm c})$;
       end for
else
       for each point $x_{j}\in\mathcal{P}_{B}$ do
             $s_{i}\leftarrow s_{i}+(\nabla\Phi)(x_{i}-x_{j})\cdot\mu_{j}$;
              // modified by $w$
       end for
end if
return $s_{i}$;
Algorithm 4 ${\rm apply\_A}(x_{i},\mu,B)$.
In Algorithm 4, $B$ is a node of the octree $\mathcal{T}$ with its $\mu$-representatives already build, and ${\rm apply\_A}(B)$ is a function that computes
$\sum_{x_{j}\ {\rm inside}\ B}(\nabla\Phi)(x_{i}-x_{j})\cdot\mu_{j}.$ (18)
To get $s_{i}$ in $s=A(\mu)$, simply call ${\rm apply\_A}(x_{i},\mathcal{T}_{\rm root})$. Note that $c$ and $w$ are hyperparameters. $c$ determines whether to use the representative as an approximation, and $w$ is the smoothing width in Sec. 4.4. Finally, we parallelize all evaluatations ${\rm apply\_A}(x_{i},\mathcal{T}_{\rm root})$ across all $x_{i}$. Other operators $A^{T}$ and $G$ are accelerated in the same way.
We remark that the idea of using a tree structure to approximate far-field interactions dates back to Barnes and Hut [4], which was originally proposed to accelerate the gravitational $N$-body problem. The tree-based algorithm is now referred to as treecode [45]. The fast multipole method (FMM) [14] pushes this idea further by using representatives for target points as well and by using expansions for propagation kernels for precision control. Previous works [3, 32] have also utilized these acceleration techniques to accelerate evaluating winding numbers. While treecodes and FMM are well-studied with publicly available libraries, to the best of our knowledge, there is currently no open-source implementation of treecode or FMM that suits our requirements: (a) GPU-based; (b) support for the operators used in our algorithm: $A,A^{T}$ and $G$. For this reason, we implement our own CUDA kernels for our algorithm.
5 Experiments
5.1 Experiment Setup
5.1.1 Implementation Details
In our implementation, the input point cloud is normalized to fit into the cube $[-1,1]^{3}$ with a margin of 1/$1/1$ 11 from the cubes boundaries. The smoothing widths are set as $w_{1}=0.002$ and $w_{2}=0.016$ unless otherwise stated. In Algorithm 4, the maximum tree depth is $D=15$, and the threshold for using an representative source is $c=2$. We run 40 iterations for each model.
5.1.2 Baselines
We choose recent state-of-the-art methods dipole propagation [34] iterative Poisson Surface Reconstruction (iPSR) [19], Parametric Gauss Reconstruction (PGR) [31] and GCNO [44] as the baselines. We use their default parameters for our experiments unless otherwise mentioned. For Dipole [34], we use orient_large for point clouds with more than 50000 points, and orient_pointcloud otherwise. All experiments are done on a desktop computer with an Intel i9-10900X CPU (3.70GHz, 10 cores) and an NVIDIA RTX 3090 GPU.
5.1.3 Datasets
Following previous work, we use datasets that are commonly used in surface reconstruction and normal orientation, including ABC [26], Thingi10k [47], the Stanford scanning repository [8, 28, 39, 13], and other datasets from Lu et al. [32], Huang et al. [20, 21], Erler et al. [12], Metzer et al. [34], Lin et al. [31], Xu et al. [44].
5.1.4 Metrics
We measure the quality of reconstructed normals with the following commonly used metrics. For point clouds with ground-truth normals $\{n_{i}^{\rm gt}\}_{i=1}^{N}$, we directly measure the angular error (AE) between reconstructed normals $\{n_{i}^{\rm recon}\}_{i=1}^{N}$ and ground-truth normals $\{n_{i}^{\rm gt}\}_{i=1}^{N}$:
${\rm AE}_{\rm pcd}=\frac{1}{N}\sum_{i=1}^{N}(1-n_{i}^{\rm gt}\cdot n_{i}^{\rm recon})/2.$ (19)
We also follow Xu et al. [44] to use the percentage of correctly oriented normals as a metric:
$P_{\rm co}=\#\{i:n_{i}^{\rm recon}\cdot n_{i}^{\rm gt}>0\}/N.$ (20)
Another common practice is to measure the accuracy of normals using their reconstructed meshes. For fairness, we use Screened Poisson Surface Reconstruction (SPSR) [23] for all methods. We set the maximum tree depth in SPSR to 10 and keep other parameters as default. We use the mesh Chamfer distance (CD) and mesh normal angular error (${\rm AE}_{\rm mesh}$). Here, ${\rm CD}$ for two meshes $M_{1},M_{2}$ is defined as
${\rm CD}=\frac{1}{2}\Big{(}\frac{1}{N}\sum_{i=1}^{N}\|x_{i}-\hat{x}_{i}\|+\frac{1}{N}\sum_{i=1}^{N}\|y_{i}-\hat{y}_{i}\|\Big{)},$ (21)
where $\{x_{i}\}_{i=1}^{N}$ uniformly samples $M_{1}$ and $\hat{x}_{i}$ is its closest point on $M_{2}$ (same for $y_{i}\in M_{2}$ and $\hat{y}_{i}\in M_{1}$). The expression for ${\rm AE}_{\rm mesh}$ is the same as ${\rm AE}_{\rm pcd}$, except it is a two-way average between the normals of the above closest samples on the meshes.
GT normalsOursiPSRDipolePGRGCNO00.5
Figure 5: Normal orientation results of different methods. We show the point clouds colored by their normals and by per-point angular error values (Eq. (19) without averaging across points). We also show the histograms of angular errors ($y$-axis in log-scale).
GT mesh and normalsOursiPSRDipolePGRGCNO
Figure 6: Normal orientation and surface reconstruction results of different methods. We show reconstructed meshes in plain colors and point clouds colored by their normals.
Table 2: Quantitative comparison results of different methods evaluated with uniform samples. The best results are marked in bold.
Method Model CD ${\rm AE}_{\rm mesh}$ ${\rm AE}_{\rm pcd}$ $P_{\rm co}$ Model CD ${\rm AE}_{\rm mesh}$ ${\rm AE}_{\rm pcd}$ $P_{\rm co}$
(×$\times 10^{-4}$) (×$\times 10^{-2}$) (×$\times 10^{-2}$) (%) (×$\times 10^{-4}$) (×$\times 10^{-2}$) (×$\times 10^{-2}$) (%)
GCNO 38.1440 11.3096 17.7151 88.4400 29.7463 12.5233 9.3252 98.3400
PGR 13.1959 11.5137 18.0785 86.8780 XYZRGB 13.5690 12.9512 21.8533 84.3820
Dipole Crab 15.1305 12.7237 12.2426 88.6625 Statuette 14.9317 13.5039 15.6297 86.2206
iPSR 2.0727 1.6770 1.9529 99.9356 2.7776 2.8389 3.4912 99.8594
Ours 1.4556 1.3515 0.8401 99.9981 2.0174 2.4180 1.7715 99.9781
GCNO 15.2068 7.2695 7.4188 99.2200 24.6961 6.7809 5.2512 99.2000
PGR 4.6810 4.3906 13.5943 93.2160 Statue of 7.4997 5.9834 11.5123 93.4080
Dipole Liberty 4.8793 4.0816 4.6036 96.2181 Delaunay 145.0533 30.4131 39.6944 60.5975
iPSR 1.5647 1.2212 1.3644 99.9963 1.5918 1.0918 1.0650 99.9963
Ours 1.0802 1.0209 0.4961 99.9975 1.1420 0.9144 0.4398 99.9994
GCNO 65.8343 44.8330 47.0361 53.9437 18.3662 8.8840 8.7805 97.6000
PGR Trash 6.4437 0.6648 2.1229 99.6740 7.9294 8.9578 19.2768 88.4840
Dipole Can 34.4190 25.5114 22.1820 77.9050 Dragon 4.0927 5.7576 5.3654 95.5606
iPSR 23.9465 14.9586 13.9820 86.4475 1.9667 5.1030 5.0781 96.3306
Ours 1.2967 0.3240 0.1370 100.0000 1.3937 4.9502 4.1447 96.3306
Table 3: Quantitative comparison results of different methods on a large-scale benchmark dataset [21]. Each metric is an average over all 1387 models used for this evaluation. The best results are in bold.
Method CD ${\rm AE}_{\rm mesh}$ ${\rm AE}_{\rm pcd}$ $P_{\rm co}$
(×$\times 10^{-4}$) (×$\times 10^{-2}$) (×$\times 10^{-2}$) (%)
Dipole 7.7909 5.3215 5.0029 95.2759
iPSR 1.9306 0.9215 0.6108 99.8208
Ours 1.1274 0.8104 0.3097 99.9328
Input pointsOursiPSRDipoleInput pointsOursPGRGCNO01
Figure 7: Qualitative comparisons for models with thin structures and sharp edges. Each model in the first two rows has 100,000 points, and each in the last two rows has 10,000 points.
[…]
Figure 8: Reconstruction results of two high-genus surfaces. We additionally show the genus $g$ of the GT models and the reconstructed meshes.
5.2 Evaluation of Normal Quality
5.2.1 Accuracy
We use uniform samples without noise to measure the accuracy of different methods. For each mesh, we use sample_surface_even in Trimesh [9] to obtain uniform samples. Each model is sampled with 160,000 points. Note that PGR [31] and GCNO [44] have prohibitively high memory and time consumption, respectively. Hence, we downsample the input point cloud to 50,000 points for PGR and 5,000 points for GCNO. A more fair comparison is provided separately in Sec. 5.2.6.
Fig. 5 shows the oriented point clouds, their angular error maps and angular error distribution. It can be seen that dipole propagation [34] has a high propability of incorrectly orienting normals in a large area. For other methods, while they can generally achieve a globally consistent orientation, their estimated normals have larger errors compared to our results. Fig. 6 shows reconstructed meshes using normals estimated by different methods. Note that almost all baselines cannot handle convex thin structures (see the interior of the trash can model). The quantitative results for these six models are reported in Table 2. Our method achieves the best overall performance.
To further validate the accuracy of our estimated normals, we conduct a large-scale study on a new surface reconstruction benchmark dataset [21]. Since we found the point clouds provided by the Huang et al. [21] contain noise (in both point positions and normals) and are unfit for quantitative evaluation, we use their provided meshes and sample uniform points using Trimesh [9]. Furthermore, some meshes provided by this dataset are non-manifold with self-intersecting components, which are not suitable for either our method or the baselines. Hence, we remove all non-manifold data and use the remaining 1387 models for this large-scale evaluation. Each model is sampled with 160,000 points. Note that PGR [31] and GCNO [44] are too computationally expensive and are thus excluded from this comparison. Table 3 reports the metrics averaged across all 1387 models. Our methods outperforms iPSR [19] and Dipole [34] in a statistical sense.
Input pointsOursiPSRDipolePGRGCNO
Figure 9: Normal estimation and reconstruction results of sparse point clouds and 3D sketches from Huang et al. [20]. Oriented points are rendered as point sprites, and we additionally render normals for 3D sketches.
5.2.2 Thin Structures and Sharp Edges
Thin structures and sharp edges constitute a difficult challenge for normal estimation methods, since normals tend to have abrupt changes around these structures. These structures also commonly exist in real-world objects, e.g., CAD models. Fig. 7 shows the angular errors for four models with thin structures or sharp edges. Our method can produce correct normals even for very thin plates, and has a lower error near sharp edges compared to other methods.
5.2.3 High-genus Surfaces
We experiment with two high-genus surfaces shown in Fig. 8. Due to their large surface areas, we sample 200,000 for each model to ensure the point clouds can depict the underlying geometric structure. We emprically found that our method and iPSR require more iterations to converge, and thus set the number of iterations for both methods to 300. In Fig. 8 we show the reconstruction results as well as the genus of each mesh. Our method can more faithfully reconstruct topological structures than other methods.
5.2.4 Sparse Points and 3D Sketches
Sparse points can be another challenge due to low sampling rates. Additionally, Huang et al. [20] studied an extreme case of sparse inputs: 3D sketches, where the points are not only sparse but also highly uniform. In Fig. 9 we show four models from Huang et al. [20]. The two uniform sparse point clouds each have 500 points, and the two sketches each have 1,000 points. Our method can obtain plausible normals for both types of point clouds.
[…]
Figure 10: Reconstruction (in yellow) and normal estimation results (in green) for a chair model with different noise levels. The normals are rendered as point sprites to indicate their orientation.
Input pointsGT meshOursiPSRDipolePGRGCNO
Figure 11: Reconstruction results for four noisy point clouds. Each model is obtained by adding $\sigma=0.5\%$ Gaussian noise to uniform samples.
Table 4: Quantitative comparisons of estimating normals for noisy models exhibited in Fig. 10 and Fig. 11. Since it does not make sense to talk about “GT normals” of noisy points, we measure only mesh errors. The best results are marked in bold.
Method Noise CD ${\rm AE}_{\rm mesh}$ Noise CD ${\rm AE}_{\rm mesh}$ Noise CD ${\rm AE}_{\rm mesh}$ Noise CD ${\rm AE}_{\rm mesh}$
level $\sigma$ (×$\times 10^{-3}$) (×$\times 10^{-1}$) level $\sigma$ (×$\times 10^{-3}$) (×$\times 10^{-1}$) level $\sigma$ (×$\times 10^{-3}$) (×$\times 10^{-1}$) level $\sigma$ (×$\times 10^{-3}$) (×$\times 10^{-1}$)
GCNO 4.1716 0.5879 2.8376 0.6831 3.9593 0.9416 7.7393 1.8608
PGR 1.5540 0.2830 1.5270 0.5409 9.3366 1.8209 6.6418 1.5744
Dipole 00 1.2309 0.5431 $0.25\%$ 2.0284 1.5253 $0.5\%$ 22.6848 5.1618 $1\%$ 14.1688 4.6009
iPSR 0.2434 0.1186 0.8082 0.3545 1.7020 0.6713 6.3857 3.1309
Ours 0.1899 0.0957 0.8242 0.3520 1.6124 0.5665 3.6541 0.8072
Method Model CD ${\rm AE}_{\rm mesh}$ Model CD ${\rm AE}_{\rm mesh}$ Model CD ${\rm AE}_{\rm mesh}$ Model CD ${\rm AE}_{\rm mesh}$
(×$\times 10^{-3}$) (×$\times 10^{-2}$) (×$\times 10^{-3}$) (×$\times 10^{-2}$) (×$\times 10^{-3}$) (×$\times 10^{-2}$) (×$\times 10^{-3}$) (×$\times 10^{-2}$)
GCNO 8.3960 9.8521 3.7648 5.4055 4.2233 11.8817 3.0652 3.2314
PGR 7.3970 10.3166 3.2521 4.8484 4.4460 18.8665 3.0096 8.6615
Dipole Cup 5.7139 23.4262 Flower 6.4231 29.9646 Buddha 5.1984 32.2373 Vase 20.2586 47.9005
iPSR 3.5387 8.5899 1.7867 4.9125 2.4111 13.7234 1.4456 3.7068
Ours 1.9513 2.9340 1.8952 4.1823 2.1011 8.8583 1.0311 2.3479
5.2.5 Noise Resilience
We study the noise resilience of different methods by adding Gaussian noise to clean uniform samples. We use three noise levels $\sigma=0.25\%,0.5\%,1\%$ (w.r.t. the bounding box diagonal) as the standard deviation for Gaussian noise. Following previous settings, each model is sampled with 160,000 points, which are downsampled to 50,000 for PGR [31] and 5,000 for GCNO [44]. We noticed that default settings of several methods do not work well with noisy inputs, and therefore we adjust the parameters to obtain optimal results. For our method, we set $w_{1}=0.02,0.03,0.04$ and $w_{2}=4w_{1}$ for the three noise levels. For iPSR [19], we set the point weight to 0.5. For PGR [31], we use their recommended settings for noisy data. Note that Dipole [34] and GCNO [44] do not have noise-adapting options or parameters, and we keep their default settings. We also change the point weight to 0.5 for the final reconstruction step using SPSR [23].
Fig. 10 shows the reconstructed meshes (in yellow) and the estimated normals (in green) for different noise levels. For dipole propagation [34], even without noise it cannot infer normals with globally consistent orientation, and as noise level increases its normals become completely misoriented. Other methods can infer a roughly correct orientation but are notably worse at preserving the geometric structure, especically at high noise levels. Fig. 11 exhibits four more noisy cases (all models are sampled with noise level $\sigma=0.5\%$). Compared with other methods, ours achieves the best balance between preserving the overall topological structure and surface details. Table 4 reports the quantitative results for the models in Fig. 10 and Fig. 11.
5.2.6 Extended Comparisons with GCNO and PGR
In our previous experiments, we have used downsampled point clouds for GCNO [44] and PGR [31] due to their high complexity. To make the comparison with PGR and GCNO more comprehensive, we rerun our method for all models used in Table 2 and Table 4 with the same number of samples as PGR and GCNO. The new metrics are reported in Table 5 and Table 6. Note that our method still outperforms these two baselines.
Fig. 12 shows the qualitative comparisons between our method, PGR and GCNO using the same number of samples. Furthermore, the authors of PGR [31] suggested that the normals solved by PGR are not really compatible with SPSR [23], and better reconstructions can only be obtained by plugging these normals back into the winding number field (WNF). While we have previously used SPSR as the reconstruction for all methods for fairness, we also test PGR with WNF-based reconstruction as in Lin et al. [31]. In Fig. 12, we observe that our method still outperforms PGR despite the final reconstruction method.
5.2.7 Real-world Scans
Different from synthetic data, real-world scans generally suffers from several kinds of imperfections, such non-uniform point distribution, scanner noise, outlier points, multi-view misalignment, occluded missing regions, etc. Furthermore, scanner data are usually large in scale, having millions of points. To evaluate the practicability of our method on scanned data, we experiment with three raw range scans (Armadillo, Bunny and Dragon) from the Stanford scanning respository [8, 28, 39, 13] and one scan (Lady) from Lu et al. [32]. Note that some raw scans contain background points as well, and we manually set a bounding box to filter out background points.
Table 5: Quantitative comparison results between our method, PGR and GCNO using the same number of samples. For comparisons with PGR and GCNO we use 50,000 points and 5,000 points for each model, respectively.
Method Model CD ${\rm AE}_{\rm mesh}$ ${\rm AE}_{\rm pcd}$ $P_{\rm co}$ Model CD ${\rm AE}_{\rm mesh}$ ${\rm AE}_{\rm pcd}$ $P_{\rm co}$
(×$\times 10^{-4}$) (×$\times 10^{-2}$) (×$\times 10^{-2}$) (%) (×$\times 10^{-4}$) (×$\times 10^{-2}$) (×$\times 10^{-2}$) (%)
GCNO 38.1440 11.3096 17.7151 88.4400 29.7463 12.5233 9.3252 98.3400
Ours Crab 20.7035 7.3027 5.7689 98.7800 XYZRGB 26.5084 12.2513 9.0382 98.0600
PGR 13.1959 11.5137 18.0785 86.8780 Statuette 13.5690 12.9512 21.8533 84.3820
Ours 3.4914 2.6714 1.8335 99.9300 5.0047 4.8284 3.4662 99.8560
GCNO 15.2068 7.2695 7.4188 99.2200 24.6961 6.7809 5.2512 99.2000
Ours Liberty 12.5331 6.8335 5.9479 99.6800 Statue of 20.6545 6.3045 4.2547 99.4400
PGR 4.6810 4.3906 13.5943 93.2160 Delaunay 7.4997 5.9834 11.5123 93.4080
Ours 3.4957 2.6623 1.2681 99.9880 3.3201 2.1595 1.0475 99.9920
GCNO 65.8343 44.8330 47.0361 53.9437 18.3662 8.8840 8.7805 97.6000
Ours Trash 102.9334 36.8877 32.5923 70.7071 Dragon 15.2683 8.2227 7.5333 98.0800
PGR Can 6.4437 0.6648 2.1229 99.6740 7.9294 8.9578 19.2768 88.4840
Ours 6.0636 0.7209 0.3859 100.0000 4.3697 5.7403 5.0017 96.3240
Table 6: Quantitative comparisons between our method, PGR and GCNO in the case of noisy inputs (same as the models used in Table. 4). For comparisons with PGR and GCNO we use 50,000 points and 5,000 points for each model, respectively. The reported metrics are for the models shown in Fig. 10 and Fig. 11. Since it does not make sense to talk about “GT normals” of noisy points, we measure only mesh errors.
Method Noise CD ${\rm AE}_{\rm mesh}$ Noise CD ${\rm AE}_{\rm mesh}$ Noise CD ${\rm AE}_{\rm mesh}$ Noise CD ${\rm AE}_{\rm mesh}$
level $\sigma$ (×$\times 10^{-3}$) (×$\times 10^{-1}$) level $\sigma$ (×$\times 10^{-3}$) (×$\times 10^{-1}$) level $\sigma$ (×$\times 10^{-3}$) (×$\times 10^{-1}$) level $\sigma$ (×$\times 10^{-3}$) (×$\times 10^{-1}$)
GCNO 4.1716 0.5879 2.8376 0.6831 3.9593 0.5879 7.7393 0.6831
Ours 0 2.5960 0.5502 0.25% 2.5552 0.6374 0.5% 3.3851 0.5502 1% 5.3490 0.6374
PGR 1.5540 0.2830 1.5270 0.5409 9.3366 0.2830 6.6418 0.5409
Ours 0.3961 0.1694 1.0456 0.3223 4.9580 0.1694 3.4486 0.3223
Method Model CD ${\rm AE}_{\rm mesh}$ Model CD ${\rm AE}_{\rm mesh}$ Model CD ${\rm AE}_{\rm mesh}$ Model CD ${\rm AE}_{\rm mesh}$
(×$\times 10^{-3}$) (×$\times 10^{-2}$) (×$\times 10^{-3}$) (×$\times 10^{-2}$) (×$\times 10^{-3}$) (×$\times 10^{-2}$) (×$\times 10^{-3}$) (×$\times 10^{-2}$)
GCNO 8.3960 9.8521 3.7648 5.4055 4.2233 11.8817 3.0652 3.2314
Ours Cup 5.0091 5.2786 Flower 3.3762 4.8993 Buddha 3.8669 10.8446 Vase 2.6241 2.6057
PGR 7.3970 10.3166 3.2521 4.8484 4.4460 18.8665 3.0096 8.6615
Ours 2.5295 2.2244 1.9771 3.8066 2.3528 8.2716 1.3408 2.3336
GT meshOursPGR normals+SPSRPGR normals+WNFGT meshOursGCNO
Figure 12: We conduct comparisons with GCNO and PGR using the same number of samples. The two models are reconstructed using 9,000 and 50,000 points, respectively. Furthermore, we additionally present the PGR reconstruction result using the winding number field (WNF) based reconstruction.
Input pointsOursiPSRDipole
Figure 13: Reconstruction results of raw range scans from the Stanford scanning repository [8, 28, 39, 13] and Lu et al. [32]. Number of points in each model: 2,374,290 (Armadillo); 362,272 (Bunny); 2,109,047 (Dragon); 530,497 (Lady). Please zoom in to see scanner noise patterns and reconstruction details.
Fig. 13 shows the reconstruction results using normals estimated by our method, iPSR [19] and Dipole [34]. To adapt to scanner noise, we set $w_{1}=0.01$ and $w_{2}=0.04$ for our method, and set the point weight to 0.5 for iPSR and the final reconstruction step. Dipole still suffers from severe misorientation (see the Armadillo head and the Dragon tail). Even though iPSR can infer a mostly correct orientation, it is less capable of handling scanner noise. Our method not only produces globally consistent orientations but also achieve a good balance between noise handling and detail preserving.
5.3 Complexity Analysis
In this section we study one important aspect of our method: the performance improvement in terms of space/time complexity. Considering the complexity of the baselines, we divide the experiments into three groups: small-scale (1,000$\sim$5,000 points), medium-scale (10,000$\sim$50,000 points) and large-scale (100,000$\sim$1,000,000 points). For the small-scale study, we experiment with all methods. We exclude PGR from large-scale experiments, and exclude GCNO from both medium-scale and large-scale experiments. For each experiment, we use a Armadillo model uniformly sampled without noise.
Furthermore, to present a more comprehensive analysis, we break down each method into a preprocessing stage and a main stage. For our method, iPSR and PGR, the preprocessing stage includes octree building, and the main stage includes normal solving. For GCNO, preprocessing includes Delaunay triangulation and the main stage includes the LBFG-S iterations. For Dipole, preprocessing incldues all steps prior to inter-patch propagation, and the main stage includes everything else. Note that the final reconstruction stage and file I/O are excluded for all methods.
In Fig. 14, we show the total running time $T_{\rm total}=T_{\rm pre}+T_{\rm main}$, preprocessing time $T_{\rm pre}$, main stage running time $T_{\rm main}$ and the total peak memory (CPU plus GPU). Time and memory measurements are shown in log-scale due to large variations between different methods. In terms of total running time, it can be seen that our method is the fastest for medium and large inputs, and even our CPU implementation (with OpenMP) is faster than other methods. Note that for small-scale inputs the CPU implementation is faster due to the overhead associated to CUDA runtime library and cross-device data transfer. In terms of memory consumption, our method is slightly larger than iPSR and GCNO for small and medium inputs, mainly because our implementation is integrated with PyTorch [2, 35] as an extension module, which involves PyTorch and CUDA overheads. However, as the input size scales up, our method has the lowest memory footprint.
Additionally, we would like to highlight our time and memory consumption improvements over the other two winding number-based methods: PGR and GCNO. Both of PGR and GCNO requires repeated evaluations of winding numbers, and both do so in a naive way by directly summing all contributions from each input point, leading to an $N^{2}$ complexity lower-bound for each evaluation ($N$ being the number of points). More specifically, to orient a point cloud with 100,000 points, PGR requires at least 3000G momery and GCNO would takes weeks. On the other hand, our treecode-based algorithm is known to have $O(N\log N)$ complexity [4]. Thus, our method easily scales up to millions of points and can finish normal orientation for these large-scale inputs in minutes.
$T_{\rm pre}+T_{\rm main}$(seconds)$T_{\rm pre}$(seconds)$T_{\rm main}$(seconds)Memory (MB)
Figure 14: Complexity comparisons of different methods, evaluated on a uniformly sampled Armadillo model with different sample numbers (first row: small-scale; second row: medium-scale, third row: large-scale). We show the total running time $T_{\rm pre}+T_{\rm main}$, preprocessing time $T_{\rm pre}$, main-stage execution time $T_{\rm main}$ and peak memory usage (MB).
5.4 Evaluation of Individual Components
In this section, we evaluate individual components of our algorithm: the smoothing width in Sec. 4.4 and the iterative steps in 4.3.
5.4.1 Smoothing width
As mentioned in previous works [32, 31] that also apply smoothing modifications to the winding number formula, this modification allows ignoring local points, leading to better robustness to noise or non-uniformity. Intuitively, a large smoothing width produces more consistent orientation but may result in oversmooth, and a small smoothing width preserves more details but may fail to orient normals consistently. In contrast to using an invariant smoothing width as in prior work [31], our width scheduling scheme that starts at a large width and ends with a small width achieves both consistent orientation and detail preserving. The first row in Fig. 15 compares using invariant smoothing widths and decreasing smoothing widths. As is evident from Fig. 15, a small width leads to orientation errors in some concave regions while a large width causes oversmoothing. Our scheduling scheme achieve consistent and detailed normal estimation.
The smoothing width can also be tuned to adjust normal smoothness at the user’s will. The second row in Fig. 15 shows three different smoothing width configurations for processing a noisy scan. A larger smoothing width leads to a smoother result.
$w_{1}=w_{2}=0.002$$w_{1}=w_{2}=0.016$$w_{1}=0.002,\ w_{2}=0.016$$w_{1}=0.01,\ w_{2}=0.04$$w_{1}=0.02,\ w_{2}=0.08$$w_{1}=0.04,\ w_{2}=0.16$
Figure 15: Study of the effect of the smoothing width. First row: decreasing the width during the iterative process allows both consistent orientation and detail preserving. Second row: adjusting widths allows explicit control over the smoothness of normals.
5.4.2 Steps in the Iterative Algorithm
We discuss the necessity of different steps in Algorithm 3. In fact, the necessity of the gradient step is easy to see: without it the algorithm simply stays at $\mu=0$ across the whole process. The rescaling step is also necessary because the scale change before and after the WNNC update is typically $10^{4}\sim 10^{10}$, depending on point density. Without rescaling the normals would soon blow up numerically.
It remains to examine the effectiveness of the WNNC updates $\mu\mapsto G(\mu)$. In Fig. 16 we show error maps of results obtained with and without the WNNC updates. These results suggest that estimated normals estimated without the WNNC updates would be highly inaccurate, especially in concave regions.
w/w/ow/w/o01
Figure 16: Ablation study of the effectiveness of WNNC updates. We show angular error maps of normals estimated with (w/) and without (w/o) using the WNNC updates. The later introduces notably larger angular error especially in concave regions.
6 Discussion
6.1 Conclusion
In this work, we propose a novel property, namely the winding number normal consistency (WNNC), derived from the winding number formula to address the problem of estimating globally consistent normal orientations. We turn the WNNC property into a practical, efficient yet embarrassingly simple iterative algorithm. We also observe the special structures in the iterative algorithm that allows acceleration using the treecode algorithm [4]. Lacking an existing package of the treecode algorithm that suits our needs (GPU-based and supporting our operators), we implement our own CUDA kernels. With extensive experiments on a variety of datasets, we exhibit the superior performance of our algorithm and implementation over recent state-of-the-art methods, in terms of both normal orientation quality and computational costs.
6.2 Limitations and Future Work
Despite the good performance of our algorithm, there are still a few aspects that could be improved or are worth exploring.
Smoothing widths
The normal estimation quality of our method still highly depends on the smoothing width. As explained in Sec. 5.4.1, an improper choice can lead to either wrong orientation or ovesmoothing. We empirically found that the widths can be kept the same for data with similar noise levels and structural complexity. Hence, adjusting the width parameters shouldn’t take much effort. However, it is still desirable for future work to explore how to automatically determine the optimal width parameters.
Non-manifold surfaces or open scans
The underlying mathematical tool of our method, the winding number formula, is meant to describe 3D shapes with a clear partitioning of space into its interior and exterior. This means our method cannot handle 3D shapes that does not satisfy this property, e.g. non-manifold surfaces and open scans. Fig. 17 shows a failure case of two planes intersecting each other. This geometry is non-manifold and does not enclose a spatial region. Our method, as well as any other implicit field-based method, is not suitable for these cases. However, as open scans are also very common in the real world, e.g., street scenes, future work should explore how to extend the winding number formula to these cases.
Further efficiency improvement
Currently we use the treecode algorithm to accelerate winding number related evaluations because it is easy to implement and parallelize. A possible future direction is to implement a more advanced acceleration technique, the fast mulipole method (FMM) [14], which further reduces complexity and allows precision control. However, FMM requires more careful mathematical analysis and is more difficult to implement and parallelize. Thus, we leave it as future work.
Figure 17: Our method is built on the winding number formula, which cannot handle non-manifold surfaces.